# Victoria, BC, Canada

### v2

@@ 3.4 Ethical Research
Ethical research is a timeless concept. However, the more recent unchecked advancements in AI and unprecedented access to publicly available datasets has meant that SE advancements are progressing faster than ethical standards, guidelines, and regulations can keep up. Both SE industry and research communities are calling for action to take stock of where we are headed collectively in terms of ethical practice in the software industry. Industrial and academic researchers alike need to be held accountable for, be aware of, and proactively manage the ethical aspects of research.
An ethical approach to research starts by carefully considering the why and not being driven by the why not. This applies to the use of any procedure, but is even more pertinent to consider with mixed methods because of the additional risks these designs bring in terms of privacy and confidentiality when mixing and integrating different types of data.
@@
  • Respect and cultural sensitivity: Many studies involve people from multiple locations and cultural backgrounds. Have we considered these cultural sensitivities when designing our protocols? Do we need to customise our protocols to make better sense to different participant groups or be more culturally appropriate in some communities? Similarly, have we considered what data sovereignty may apply to special groups such as Indigenous peoples [^6] https://static1.squarespace.com/static/5b3043afb40b9d20411f3512/t/63ed934fe861fa061ebb9202/ 1676514134724/Communique-Indigenous-Data-Sovereignty-Summit.pdf?
@@ 4 A Landscape of Mixed Methods Design
Research study design that involves choosing suitable procedures and deciding on their timing is a challenging process easterbrook_selecting_2008 , but research design is particularly challenging in the case of an MMR study. The research problem or questions being asked and how they emerge over the course of a study guides the research design and the timing of procedures used. Below we discuss the main characteristics of a research design, followed by some common patterns of MMR designs.
4.1 Core Characteristics of Mixed Research Designs
  • Planned or Emergent Design: A mixed research design may be planned up front as guided by the research problem or question, or it may organically evolve over time because a first approach may not work as well as anticipated (for example, due to low participation), or it may lead to insights that beg for confirmation, expansion, explanation, or triangulation.
@@ 4.1 Core Characteristics of Mixed Research Designs
  • Inductive/Deductive Dominance: An MMR design typically involves both inductive (from data to theory) and deductive research (from theory to data), but one approach usually dominates or takes priority over the other driven by the philosophical world view of the researcher or by the nature of the research problem or question. Researchers often conflate inductive and deductive research with quaL and quaN methods, respectively, but Young et al. explain how inductive and deductive research may fall along a “spectrum” of quaL and quaN data young_spectrum_2020 .
  • Timing of Procedures: The various procedures may be conducted in sequence or in parallel. A sequential design is driven by the emergence of new or refined research questions, while a parallel approach may be motivated by the expertise of the research team, access to research participants and data, or timing constraints.
  • Data Integration: How the data is integrated during collection or analysis shapes the research design and the research insights ladner_mixed_2019 (as mentioned above). For example, in the case of parallel MMR, quaN and quaL data may be merged to lead to the findings or insights, or in the case of a sequential study, one type of data may build on top of or add to another type (see Figure 1).
@@ 4.2 Research Designs
The characteristics mentioned above and how they are composed lead to a constellation of different research designs. We use the nomenclature for research designs as proposed by Creswell and Plano Clark creswell_best_2011 while describing some popular design variations in the following. As we describe each design variation, we provide an example of a recent paper published in a relevant empirical software engineering journal that uses that design. [^7] We found looking through the last twelve issues of the Empirical Software Engineering journal and the Information and Software Technology journal, that many SE papers use mixed methods, but they often do not say they are mixed methods research, and they also usually do not articulate or justify the specific design they use. Here, we provide one example paper that maps to each research design. If the paper was not that clear about the design, we contacted the authors to confirm our mapping.
@@
  • Exploratory Sequential: This sequential design often prioritizes a quaL method (normally the dominant method) followed by a quaN procedure that builds on insights from the first procedure. The design may be planned up front or it may emerge as insights are uncovered from the first procedure. For example, Alami et al. Alami2024 explore how psychological safety impacts agile team behavior through a mixed method approach consisting of an initial exploratory phase with 20 interviews, followed by a quaN survey study involving 423 participants.
@@
  • Explanatory Sequential: This sequential design starts with and prioritizes a deductive quaN method (normally the dominant method) that is followed by a quaL method to help explain the insights from the first procedure. The design may be planned up front or the need for the second procedure may emerge, especially in light of some unexpected insights or where some insights need further interpretation or explanation. For example, Jiang et al. Jiang2024 mined open source deep learning projects to analyze the distribution of computer vision defects, followed by interviews conducted with leaders of a reengineering team to shed light on the challenges and practices that may have led to the defects.
@@
  • Convergent Parallel: The convergent parallel design involves both a quaL procedure and a quaN procedure to bring complementary insights on a research question. The goal may be to uncover both convergent and divergent insights from both procedures that are planned in parallel. For example, Hidellaarachchi et al. Hidellaarachchi2024 carry out personality test-based surveys and in-depth interviews with a selected set of participants to investigate the impact of personality on requirements engineering activities. The survey and interviews were conducted in parallel and insights obtained from both approaches were used to answer the research questions.
  • Embedded: An embedded or nested design is when one procedure is dominant (quaL or quaN) but another procedure appears as a “strand” within the first procedure. For example, as part of a quaN survey, a researcher may choose to interview select participants while carrying out the survey to gain more quaL insights on their answers. A representative paper that describes using this research design can be found in the work by Madugalla et al. Madugalla2024 where the mixed (quaL+quaN) survey procedure is dominant, followed by nested (quaL) interviews. Their study aimed to identify challenges and possible benefits of conducting SE research with human participants during the pandemic. It nested the findings of mixed survey data with in-depth follow up interviews with those researchers who had answered the survey and volunteered for the follow up interviews.
@@ 5 MMR Design Scenarios
To illustrate the principles, characteristics and MMR designs we presented so far in this paper, we showcase four scenarios as guiding examples. There could be many more study design examples, but we focus on four that map to each of four MMR designs (exploratory sequential, explanatory sequential, convergent parallel, and embedded) for the sake of brevity and clarity. For each scenario, we suggest a fictional SE researcher who is pursuing a mixed method design, and use these characters to show how having empathy and understanding for our study participants can also lead to more thoughtful and rigorous research.
We describe the four scenarios below and mention the decisions and challenges each fictional researcher could have faced in designing and conducting their research. For the four scenarios, we provide illustrations of the research designs (see Figures 3, 4, 5 and 6). We also summarize the four scenarios in Table 1.
Ali (he, him), is a postdoc working closely with a partner company studying the introduction of a new security tool that is used by development teams. The use of the tool enforces the compliance of security practices by developers at this large company. Introducing a new tool like this that enforces changes in development work has complex impacts not just on the nature of work, but also on the developers’ lived experience. Ali’s research question is to investigate how this tool impacts perceived developer productivity as well as their experience while developing. Ali has prior industry experience and recognizes the complexity of introducing this new tool and decides to use mixed methods to bring and integrate insights on the technical and social aspects of the new tool.
As Ali finds little related research to build on, he first conducts an inductive exploratory open-ended interview with a small number of developers from a team at his partner (case) company to find out how this change has impacted their developer practices for writing more secure code, and also how it has impacted their perceived productivity and experience. As Ali conducts these interviews, he takes care to protect the identity of the team members he interviewed, with both the security of the organization and ethical concerns with respect to participant job security in mind. For example, developers may be frustrated by such a tool but may not want to openly share this at their company. The findings from these interviews helps identify a number of concepts that reveal changes in developer productivity and experience, in addition to changes in the perceived quality and security of the code.
Ali finds that, contrary to his expectations from his prior industrial experience with mandated security tools, the tool does not negatively impact perceived developer productivity nor does it negatively impact their experience. He wishes to know if the interview findings from the members on this team hold (generalize) across a broader set of developers at his partner company. He also learned that some developers changed the tools they were using, a finding he did not expect. He expands the preliminary exploratory approach and uses the insights from the initial interviews to develop a follow-up closed question survey with a larger population of developers at the same company. In addition to asking questions about their perceived productivity, and experience, he also asks additional questions about the tools they use, an expansion over what he intended originally. In the survey, the identity of the developers and their projects is again protected.
Both methods were conducted in a rigorous way and led to novel insights about the impact of a security compliance tool on perceived developer productivity and their experience. Through the survey he also learned answered some questions prompted by the interviews across a larger population of developers. Ali wanted to conduct both procedures (sequentially) at the start of the research and considers the qualitative insights as the dominant contribution in this exploratory sequential research design.
The trade-off Ali makes in this design is to focus some of his research effort on the follow-up survey to triangulate, to generalize, the findings over achieving potentially deeper insights from longer or more interviews in the first phase. But the benefits of using both approaches is that he gains insightful results and finds they hold across more teams and in different team contexts. See Figure 3 for an overview of Ali’s research approach.
Sam (they, them) is an early stage Ph.D. student collaborating with a large company that is trialing a new Generative AI tool to automate part of the code review process with hypothesized improvements on time-to-merge of code changes and on the quality of the reviewed code.
Sam starts their research with a deductive quantitative data mining procedure to investigate their hypothesis that this new tool (before and after use) improves time-to-merge and reduces the number of bugs captured in their industry partner’s code repository. As they mine the data, Sam ensures that their findings do not reveal whose coding or code review activities may have resulted in the future introduction of bugs with other team members or leads.
From the mining of data, Sam confirms the tools lead to the hypothesized faster time-to-merge but not a lower number of bugs after merge. Sam wishes to understand this unexpected result, and so to gather the views of the developers who used this tool, Sam conducts open ended interviews with a random selection of developers to probe inductively why it did not have the hypothesized impact on bug count.
Care is taken to protect the identity of the developers interviewed, but also to ensure the interviews did not make participants feel stressed given the presence of some bugs they may have introduced. From the interviews, as well as the mined data, novel insights about the tool and how it was used, emerged. Notably, from the interviews, Sam learns that the developers are able to work more efficiently but they are perhaps a little more complacent using the AI reviewing tool.
The use of mixed methods was not intended up front, but was an emergent or organic research design decision so that an explanation could be found through the interviews. The integrated insights can be used to shape code review tool design accordingly, by adding features to prompt developers to avoid complacency.
The resulting research design Sam followed is an explanatory sequential design (see Figure 4). A trade-off was made to halt further mining of the telemetry data indicating speed of merge and number of bugs, to instead conduct interviews with developers to explain the surprising results, but also to shed light on more actionable insights for how the code review tool could be improved.
Vicki (she, her) is an experienced researcher that has been studying collaboration practices in open source projects for some time. Her most recent research aim is to understand the adoption of a new onboarding process to address the retention of new contributors to an open source project that she has been involved with (as a contributor and researcher) for some time.
The onboarding process relies on a form of mentoring that has not been studied before. This process suggests that seasoned and new developers pair on work items, asynchronously communicating over Slack).
She decides to inductively consider the onboarding process from two angles in parallel. Vicki uses observations of their communication over the project’s asynchronous communication channel to discern how the onboarding process is used by both existing and new contributors in the project. See Figure 5 for an overview of her research design. Note, because of ethical concerns, Vicki anonymizes the identity of the participants she observes.
In parallel to, and to complement her observations, she mines logged data about code commits from the new contributions that are made by the new contributors. She studies the quality of their first commits but also considers the timing and quality of their subsequent commits.
She recognizes that mining the commit data comes with the potential risk of introducing privacy concerns, so she obfuscates developers’ identities. She finds that mentoring during the onboarding process leads to higher quality commits and an increase in commits for those who were mentored.
The use of a convergent parallel design that involves conducting both qualitative and quantitative procedures in parallel provides complementary and novel insights on the adoption of this new onboarding process in this open source community. She finds that the onboarding mentorship process of pairing and communicating over Slack leads to an increase in future contributions and higher quality commits.
Zara (she, her) is a late stage Ph.D. student who aims to investigate how effective an automated tool she designed is for generating tests, in comparison to not using the tool. She initially conducts a lab experiment with senior student developers, where half use the tool and half do not. She takes care during recruitment that the students’ professor is not involved and thus may coerce students to participate (an ethical concern).
She starts, deductively, with the hypothesis that the tool will lead to faster test generation. She assigns the same testing tasks to both groups. Zara finds that the student developers are faster at writing tests using the automated tool. For a random selection of the participants, thinking inductively, she asks the students to qualitatively describe their confidence about the tests they created, and how the tool impacts their knowledge about testing in general, through intermittent short surveys they fill out between tasks. She finds from the two methods, that the students using the automated tool are not only faster, but they also report more confidence in their tests and learned more about testing.
An alternative design to the mixed methods design could have been to mine data from the students’ activities as a single method. However, she used two mixed methods as she was also interested to gain insights on how the tool impacted their confidence and learning about testing. Another advantage of using the embedded surveys, is that the surveys also helped explain, and lead credibility to, the quantitative results from the mined data.
The embedded design she used is primarily a quantitative method, examining the speed of tests created, but has an embedded qualitative component, to find out additional insights about their perceived confidence and learning (see Figure 6 for an overview of this MMR design).
Type of Design Design flow Rationale(s) Scenario
@@ 5.4 Zara’s Study: Evaluating the Effectiveness of an Automated tool for Generating Tests
quaL: Inductive Open-ended Interviews $\rightarrow$
quaN: Deductive Closed Survey
@@ 1. Ali
2. Sam
AI Code review
@@ 2. Sam
quaL: Inductive Ethnographic Observations $||$
@@
3. Vicki
Adoption of
Onboarding Mentorship
@@ 3. Vicki
quaN: Deductive Lab Experiment $\rightarrow$
quaL: Surveys
@@ 4. Zara
Table 1: Four fictional research scenarios demonstrating four different possible designs for mixed method research.

### v3

@@ 3.4 Ethical Research
Ethical research is a timeless concept, with a push by institutions and publishing organizations towards the use of Ethics’ Boards to review studies for ethical compliance. However, the more recent unchecked advancements in AI and unprecedented access to publicly available datasets has meant that SE advancements are progressing faster than ethical standards, principles, and regulations can keep up. Both SE industry and research communities are calling for action to take stock of where we are headed collectively in terms of ethical practice in the software industry. Industrial and academic researchers alike need to be held accountable for, be aware of, and proactively manage the ethical aspects of research.
An ethical approach to research starts by carefully considering the why and not being driven by the why not. This applies to the use of any method, but is even more pertinent to consider with mixed methods because of the additional risks these designs bring in terms of privacy and confidentiality when mixing and integrating different types of data.
@@
  • Respect and cultural sensitivity: Many studies involve people from multiple locations and cultural backgrounds. Have we considered these cultural sensitivities when designing our protocols? Do we need to customise our protocols to make better sense to different participant groups or be more culturally appropriate in some communities?
@@ 4 A Landscape of Mixed Methods Design
Research study design that involves choosing suitable methods and deciding on their timing is a challenging process easterbrook_selecting_2008 , but research design is particularly challenging in the case of an MMR study. The research problem or questions being asked and how they emerge over the course of a study guides the research design and the timing of methods used. Below we discuss the main properties of research design, followed by different types of integration, and common patterns of MMR designs.
4.1 Design Properties
  • Research Questions: MMR questions differ from purely quantitative questions (that typically focus on description, comparison, or relationships between variables) or qualitative questions (that are typically open-ended, evolving, and exploratory witgh the aim to uncover processes or experiences). The MMR question addresses what the researcher hopes to learn through the combination (integration) of the quaL and quaN data creswell_designing_2018 . In turn, these questions will shape other design properties as listed next.
  • Planned or Emergent: An MMR design can be either planned (i.e., predetermined), where it is structured upfront based on the research problem or question, or emergent, where it evolves organically throughout the study. The emergent approach may occur if the initial method proves ineffective (e.g., due to low participation) or if initial findings generate new insights that require confirmation, expansion, explanation, or triangulation.
@@ 4.1 Design Properties
  • Inductive/Deductive Dominance: An MMR design typically involves both inductive (from data to theory) and deductive research (from theory to data), but one approach usually dominates or takes priority over the other driven by the philosophical world view of the researcher or by the nature of the research problem or question. Researchers often conflate inductive and deductive research with quaL and quaN methods, respectively, but Young et al. young_spectrum_2020 explain how inductive and deductive research may fall along a “spectrum” of quaL and quaN data.
  • Timing: Time orientation in MMR design refers to the type of implementation process, specifically focusing on when quaL and quaN data collection occurs. This can happen sequentially (i.e., conducting first one quaL/quaN method and then the other quaL/quaN) or in parallel (i.e., conducting the two quaL/quaN methods at the same time). For some cases, a sequential design is driven by the emergence of new or refined research questions, while a parallel approach may be motivated by the expertise of the research team, access to research participants and data, or timing constraints.
Additional design properties could be included in this list, such as epistemological perspectives and paradigmatic assumptions from a decision-making viewpoint, or strategies like sampling, data collection and data analysis approaches (as also discussed in Venkatesh et al. venkatesh2016guidelines ). However, for the sake of simplicity, we focus on these key aspects: Research Questions, Planned or Emergent, Inductive/Deductive Dominance and Timing, complemented by a discussion on the Types of Integration in the following subsection. Nevertheless, we acknowledge that all these design properties require careful consideration by researchers, as they significantly influence the choice of MMR design type (explained further in Subsection 4.3).
4.2 Types of Integration
Integration serves as a foundational element of MMR, as noted in Section 2 and depicted in Figure 1, and should be thoroughly detailed reporting any MMR study. Creswell and Plano Clark endorse this view (poth_sage_2023, , Chapter 2) and explain that a comprehensive description of integration, considering both its intent and procedural aspects, enhances transparency by demonstrating how different methods were effectively employed. Below we present four types of integration in MMR, as recommended by Kuckartz and Rädiker (poth_sage_2023, , Chapter 22).
  • Sequential integration occurs when the findings from one method inform the application of another method. For instance, the results of an initial quaN method may guide the development of quaL data collection instruments or inform sampling strategies for the following phase.
  • Results-based integration involves combining, comparing, and relating findings only after the analyses of both quaL and quaN methods have been completed. This type of integration focuses on synthesizing the outcomes of each method (quaL and quaN) to derive comprehensive insights.
  • Data-based integration involves the simultaneous analysis and interrelation of quaL and quaN data during the analysis phase. This type of integration requires that both data types are associated with the same underlying source, ensuring that quaL and quaN data can be distinctly assigned and connected to each individual case. This alignment allows for a more cohesive and in-depth examination of the research subject.
  • Transformation-based integration occurs when quaL data is converted into quaN data, or vice versa, allowing for integration within a single “data world” (quaL or quaN). This type of integration enables researchers to analyze transformed data using the corresponding analytical techniques, enhancing the depth and breadth of the analysis.
Additional examples of integration can be found in Onwuegbuzie and Johnson’s Guide to Mixed Methods Analysis onwuegbuzie_routledge_2021 and Poth’s book poth_sage_2023 —the latter includes a dedicated section on “designing innovative integrations with technology” organized by Guetterman, who introduced the idea of joint displays during integration in MMR (see guetterman_integrating_2015 ; guetterman_visuals_2021 ). Di Penta and Tamburri Dipenta also propose the adoption of mind maps to gain an overview of the design to ensure there is alignment with the study’s quaN/quaL focus, and to aid in identifying gaps for further exploration, and to address threats to validity. Venkatesh et al. venkatesh2016guidelines provide more detailed prescriptive guidelines to assist in the integration and evaluation of MMR for Information Systems research, some of which may be useful for SE research.
@@ 4.3 Research Designs
The design properties mentioned above and how they are composed lead to a constellation of different research designs. We use the nomenclature for research designs as proposed by Creswell et al. creswell_best_2011 while describing some popular design variations in the following. As we describe each design variation, we provide an example of a recent paper published in a relevant empirical software engineering journal that uses that design. [^9] We found looking through the last twelve issues of the Empirical Software Engineering journal and the Information and Software Technology journal, that many SE papers use mixed methods, but they often do not say they are mixed methods research, and they also usually do not articulate or justify the specific design they use. Here, we provide one example paper that maps to each research design. If the paper was not that clear about the design, we contacted the authors to confirm our mapping.
@@
  • Exploratory Sequential: This sequential design often prioritizes a quaL method (normally the dominant method) followed by a quaN method that builds on insights from the first method. The design may be planned up front or it may emerge as insights are uncovered from the first method. For example, Alami et al. Alami2024 explore how psychological safety impacts agile team behavior through a mixed method approach consisting of an initial exploratory phase with 20 interviews, followed by a quaN survey study involving 423 participants.
@@
  • Explanatory Sequential: This sequential design starts with and prioritizes a deductive quaN method (normally the dominant method) that is followed by a quaL method to help explain the insights from the first method. The design may be planned up front or the need for the second method may emerge, especially in light of some unexpected insights or where some insights need further interpretation or explanation. For example, Jiang et al. Jiang2024 mined open source deep learning projects to analyze the distribution of computer vision defects, followed by interviews conducted with leaders of a reengineering team to shed light on the challenges and practices that may have led to the defects.
@@
  • Convergent Parallel: The convergent parallel design involves both a quaL method and a quaN method to bring complementary insights on a research question. The goal may be to uncover both convergent and divergent insights from both methods that are planned in parallel. For example, Hidellaarachchi et al. Hidellaarachchi2024 carry out personality test-based surveys and in-depth interviews with a selected set of participants to investigate the impact of personality on requirements engineering activities. The survey and interviews were conducted in parallel and insights obtained from both approaches were used to answer the research questions.
  • Embedded: An embedded design nests a secondary method within a primary method with the intent up front to integrate the data from both methods to arrive at the study findings (that is, it relies on data-based integration as opposed to results-based integration we may see in a sequential design). In the study by Madugalla et al. Madugalla2024 which aimed to identify challenges and possible benefits of conducting SE research with human participants during the pandemic, research findings emerged from the integration of the data obtained with a survey (quaN and dominant method) with data from the interviews conducted with a sample of the same participants (quaL and nested method). Yu and Khazanchi yu2017 offer practical advice for conducting embedded MMR designs.
@@ 5 MMR Design Scenarios
To illustrate the principles, properties and MMR designs we presented so far in this paper, we showcase four scenarios as guiding examples. There could be many more study design examples, but we focus on four that map to each of four MMR designs (exploratory sequential, explanatory sequential, convergent parallel, and embedded). For each scenario, we suggest a fictional SE researcher who is pursuing a mixed method design, and use these characters to show how having empathy and understanding for our study participants can also lead to more thoughtful and rigorous research. We describe the four scenarios below and mention the decisions and challenges each fictional researcher could have faced in designing and conducting their research in terms of the four principles we summarize in the paper. For the four scenarios, we provide illustrations of the research designs (see Figures 3, 4, 5 and 6). We also summarize the four scenarios in Table 1. The study scenarios do not mention all the necessary details expected in a paper, as we focus on illustrating the principles and properties of mixed method research in SE to match the goal of this paper.
Scenario Type of Design Design Flow Rationale(s) Integration
5.1 Ali
Security Compliance
Tool
@@ 5.1 Ali
quaL: Inductive Interviews $\rightarrow$
quaN: Deductive Survey
@@ 5.2 Sam
Sequential
5.3 Vicki
Adoption of
Onboarding Mentorship
@@ 5.3 Vicki
quaL: Inductive Observations $||$
@@
Results-based
5.4 Zara
Test generation
@@ 5.4 Zara
quaN: Deductive Lab Experiment $||$
quaL: Deductive Surveys
@@
Table 1: Four fictional research scenarios demonstrating four different possible designs for mixed method research, showing the flow of the methods, the rationale(s) and type of integration for each design.
Ali (he, him), is a postdoc working closely with a partner company studying the introduction of a new security tool to ensure compliance with secure development practices across a large software company. Ali and his partners recognize that introducing a tool that enforces changes in development work has complex and likely unexpected impacts, not just on the security and or quality of the authored code, but also on the developers’ perception of their productivity and experience. His research goal is to identify what impacts the tool may have (within the context of the developers’ work and environment), and using these initial insights, investigate if these perceptions are shared broadly by other developers across the company. Ali’s research question asks: How does a development tool that enforces compliance with secure development practices impact developer perceived productivity and their experience?
Methodological Rationale:
Through a literature review, Ali finds a lack of research on developers’ perceived impacts of the specific type of security tool he is studying. He aims to investigate whether developers believe that the tool improves the quality of the code but also how it may impact other aspects of their development experience. Ali could begin by designing a mono-method survey, such as one based on a validated technology adoption model tam89 , and distributing it widely across the company. However, his primary goal is to explore inductively what impacts the new security compliance tool might have, focusing on both the anticipated and unanticipated aspects of developers’ experiences. Furthermore, Ali aims to assess whether these initial findings can be generalized to a broader population of developers within the company. Ali’s goals could not be achieved without an integrative thinking approach supported by MMR, which leads him to use an exploratory sequential mixed methods design.
First, Ali interviews (quaL method) about 20 developers to uncover these impacts. Next, based on the initial insights, he conducts a follow-up survey (quaN method) across a broader population of developers at the studied company. A survey that is faster to answer was part of his rationale for this design. See Figure 3 for an overview of Ali’s MMR approach.
Novelty of Integrated Insights:
The sequential integration of findings from both methods provided novel insights into the impact of a new security compliance tool on developers’ perceived productivity, workflow, and learning experiences. Furthermore, these impacts were found to be generalizable to other developers. Through interviews, Ali finds the tool does not negatively impact perceived developer productivity in terms of velocity, and that it surprisingly increases the developer’s flow experience and helps the developers learn about and appreciate new security practices. He also uncovers that some developers abandoned using tools they were fomerly using because of this new tool, another finding he did not expect. These anticipated and unanticipated findings were used by Ali to develop his follow-up survey with closed survey questions to investigate quantitatively if these experiences generalized across a broader set of developers at his partner company.
Procedural Rigor:
The interviews and survey were conducted following respective best practices [^10] We do not elaborate on the rigor of the individual methods used to focus our paper on the discussion of rigor that applies to the mixing of methods.. As mentioned above, the rationale for using these two methods was primarily to develop the survey and to triangulate and generalize the findings across a broader sample. Ali aimed to conduct both methods (sequentially) at the start of the research, but considers the qualitative insights as the dominant contribution in this exploratory sequential research design. Ali, was careful not to invite those interviewed to answer the survey, as doing so could influence their survey responses. Ali also took care to word survey questions to avoid bias with insights gained through the interviews. Ali assesses the overall findings from MMR, not from individual studies but as an integrated contribution. He discusses the potential threats to validity that may arise during data collection and analysis when compiling the final results report for his study. For instance, it could have been the case that the follow up survey led to divergent results about the abandonment of other tools or on productivity experiences. Such divergences would need further investigation.
Ethical Considerations:
As Ali conducts both methods, he takes care to protect the identity of the team members he interviewed and surveyed. Ensuring this confidentiality was especially important given security is a key concern of the organization, and developers may not feel free to share concerns about the friction a compliance tool may introduce. Ali also was careful how nuanced insights from the interviews (e.g., about abandoned tools which may have been used by specific developers only) were later framed in a neutral ways so that the small group of interviewed developers could not be identified. Ali also took care not to send the survey to the interviewed participants, both to ensure the integrity of the data (supporting rigour), but also to be respectful of participant time. The design itself was also done with the careful use of developer time in mind.
Sam (they, them) is an early stage Ph.D. student collaborating with a large company that is trialing a new Generative AI tool to automate part of the code review process. Sam’s initial research goal is to test their hypothesis that a new GenAI tool that automates the code review process will improve the time to merge of code changes and improve the quality of the reviewed code (by reducing the number of bugs in the newly committed code).
Methodological Rationale:
Sam starts their research with a deductive data mining method (quaN method) to investigate their hypothesis that the GenAI reviewing tool (before and after use) improves time-to-merge and reduces the number of bugs captured in their industry partner’s code repository. They do not start with the intention to follow a mixed methods research design, and were planning to conduct the data mining study only. However, from the mined data Sam confirms the tool leads to the hypothesized faster time-to-merge but not a lower number of bugs after merge. Sam wishes to understand this unexpected result. To gather the views of the developers who used this tool, Sam feels compelled to conduct open ended interviews (quaL method) sequentially sampling a subset of the developers to probe inductively why the tool did not have the hypothesized impact on bug count. The resulting research design Sam followed is an explanatory sequential design (see Figure 4).
Novelty of Integrated Insights:
From the interviews combined with the mined data, novel insights about the tool’s efficacy and how it was used, emerged. Notably, from the sequential integration of the mined data and interviews, Sam learns that the developers are able to work more efficiently but they are potentially more complacent when using the AI reviewing tool. Sam reflects on the mixed insights and their implications, limitations, and future directions. For instance, these integrated insights can be used to shape code review tool design accordingly, by adding features to prompt developers to avoid complacency.
Procedural Rigor:
A mixed methods design was not intended up front, but was an emergent or organic research design decision to explain the surprising results from the data mining. In addition to ensure the mining process and interviews were conducted with rigor, Sam must ensure that the sequential integration step was also done in a way to reliably explain the surprising result from the mined data. As Sam may be biased following this unexpected finding, they recruit another student to conduct these interviews. Note Sam, could have used an embedded design with a minor survey strand to explain the insights from the mined data but the findings were so unexpected they anticipated that interviews could also inform how the review tool could be improved before more mining.
Ethical Considerations:
As Sam mines the telemetry data, they ensure that their findings do not reveal whose coding or code review activities may have resulted in the future introduction of bugs with any other team members or managers, as this could be used to evaluate their performance. Care is also taken by Sam to protect the identity of the developers interviewed, and also to ensure the interviews did not make participants feel stressed given the presence of some bugs they may have introduced that was found in the telemetry data and that they feel could be linked to them directly.
Vicki (she, her) is an experienced researcher that has been studying collaboration practices in open source projects. Her most recent research goal is to understand the adoption of a new onboarding process to address the retention of new contributors to an open source project that she has been involved with (as a contributor and researcher) for several months.
Methodological Rationale:
The onboarding process relies on a style of mentoring that has not been studied before and suggests that seasoned and new developers loosely pair on work items, asynchronously communicating over Slack. As this is a new process, Vicki decides to inductively consider the onboarding process from two angles in parallel. Vicki uses observations (quaL method) of participant communication over the project’s asynchronous communication channel to discern how the onboarding process of pairing is used by both existing and new contributors in the project. In parallel to, and to complement her observations, she mines logged data (quaN method) about code commits from the contributions that are made by the new contributors. She studies the frequency and timing of commits during and following the informal pairing activities over Slack. See Figure 5 for an overview of her MMR design.
Novelty of Integrated Insights:
Vicki first analyzes the qualitative and quantitative data to understand the nature of the communication and the impact of onboarding on commits. She then conducts a results-based integration by combining, comparing and relating findings from the qualitative communication data and the quantitative commit data. This integration leads to the meta-inferences that contribute theoretical insights about how pairing during the onboarding process may lead to faster commits. The use of a convergent parallel design that involves conducting both quaL and quaN methods in parallel provides complementary and novel insights on the adoption of this new onboarding process in this open source community. Vicki also finds that the onboarding mentorship process of pairing and communicating over Slack leads to an increase in future contributions and higher quality commits.
Procedural Rigor:
The choice of a convergent parallel design supports Vicki to inductively analyze the communication about pairing qualitatively and link the communication instances to the commits that are potentially happening in parallel (or soon after). This design allows her to integrate and build a story about how the mentors and new contributors communicate and pair, and the impact of that communication on the timing and presence of subsequent or in parallel commits. She considers the quantitative aspects (number and timing) of the commit data in an inductive way to see what emerges (she has no initial hypothesis about the impact of this informal pairing over Slack). As she integrates the two data sets, she is careful to specify the timing intervals as she links the two sets of data (the timing interval used is suggested by other research that links these two types of data). Together, these two data sources lead to complementary insights and theoretical insights about the nature of the communication about pairing and the commits the communication is connected to. Her analysis leads to meta-inferences about the pairing activities and commit behaviour, notably that pairing during onboarding lead to faster commits. The rigor of her design is increased by checking the coding reliability with other researchers and running statistical analysis on the quantitative data she collects.
Ethical Considerations:
Vicki seeks permission from her research participants and project owners to both monitor the project communication channels and mine logged data about code commits and their comments. She also ensures that new contributors agree to having their communications with the mentors analyzed for the study, and that they may opt out at any time (ensuring any collected information is deleted from the study data across both methods). Likewise, mentors also must give permission and can opt out at any time. She anonymizes the identity of the participants she observes in the communication channel, ensuring no identifying data remains in the reported data. Additionally, during the results presentation, she aggregates the participant’s profile to avoid putting any particular group in evidence, as this was not part of her study’s goal (e.g., gathering only relevant personal data for her research and not showing personal characteristics in her participant’s demographic table). Vicki recognizes that mining the commit data comes with the potential risk of introducing significant privacy concerns, so she also obfuscates the commit data. Vicki conducts her observations of the Slack channels in a transparent but nonobtrusive manner, ensuring security in its storage and planned deletion after the study.
Zara (she, her) is a Ph.D. student who aims to investigate how effective an automated tool she designed is for generating tests. Zara faces the challenge of evaluating a tool prototype she developed without an industry partner, and she does not have access to developers in industry for evaluating her tool. Her research hypothesis is that her tool will lead to faster test generation when used by senior computer science students.
Methodological Rationale:
Zara starts, deductively, with the hypothesis that use of her tool will lead to faster test generation over standard test generation features available in the IDE. Zara decides to conduct a lab experiment (quaN method) with senior student developers who are well versed in software testing techniques, where half of the students she recruits use the tool and the other half do not. Zara also anticipates that the students may have more confidence about the tests they authored and may learn about testing through use of her tool. Inviting all participants, she decides to run an embedded data collection method, where she asks them to describe their confidence about the tests they create, and how the tool or IDE they are using impacts their knowledge about testing through intermittent short surveys (quaL method) presented between tasks. Without this embedded qualitative method, she would not gain insights into their confidence of their tests or how they learned more about testing. The embedded deductive design she used is primarily a quantitative method, examining the speed of tests created, but has a smaller embedded deductive qualitative component, to report on their perceived higher levels of confidence and learning. Note that the secondary method relies on the dominant method leading to the embedded design decision. See Figure 6 for an overview of this MMR design.
Novelty of Integrated Insights:
Zara’s study reveals that using the automated tool for test generation not only improves the speed of test creation compared to standard IDE features (identified from the experiment) but through the embedded survey she finds it also enhances the participants’ confidence in the tests they generate and their overall understanding of software testing concepts. Using a data-based integration that involes the simultaneous analysis of both quantitatve and qualitate data brings this new meta-inference about the impact of this new tool (when compared to the features in an IDE) on student perceived confidence and learning.
Procedural Rigor:
As Zara designed the study, Zara invited two other graduate students in her lab to validate the data collection instrument developed/used. Participants are selected randomly for each condition, and those who are not assigned to use her tool are free to use other test generation features available in modern IDEs. Zara assigns the same testing tasks to both groups in this between groups design, and asks similar questions to both groups in the embedded survey to tease out how the used tools impacted their confidence in testing and their learning of testing concepts. The advantage of using the embedded survey, is that it helped explain, and lead credibility to the quantitative results from the mined data.
Ethical Considerations:
The use of student participants calls for several ethical considerations. Participation is voluntary and their identity is kept confidential across both methods. Zara takes care during recruitment that the students’ teachers are not involved in the study or research, that the teachers do not know who participates in the study, and that they do not put any real or apparent pressure on students to participate and thus don’t introduce power dynamics. The students are reassured that their decision to participate (or not) and nature of the feedback they provide will have no bearing on their performance in the class. The students’ time to participate is also valued. The potential benefit of gaining experience with test generation is communicated to the students during recruitment (this value was validated with teachers before the study). Care is taken to ensure students do not feel stressed during the study, and they can opt out at any time.